OpenStudy (debpriya):

A trader bought an article and then sold. Had he bought at 10% less and sold at 10% more, he would have doubled his profit percentage. What percentage of the selling price is the cost price ?

OpenStudy (mathstudent55):

He bought it for x and sold it for y. The profit percentage is \(\dfrac{y - x}{x} \times 100\) The profit percentage could have been \(\dfrac{1.1y - 0.9x}{0.9x} \times 100\) The second percentage is twice the first percentage, so you have \(\dfrac{1.1y - 0.9x}{0.9x} \times 100 = 2 \times \dfrac{y - x}{x} \times 100\) Solve for x in terms of y.

OpenStudy (debpriya):

Thank you :)

OpenStudy (mathstudent55):

You're welcome.